Two resistors, 38.0 Ω and 40.0 Ω, are connected in parallel. The current through the second resistor is 3.00 A.


(a) Determine the current in the first resistor?

(b) What is the total power supplied to the two resistors?
Answer #1

Voltage across Parallel resistance is same-
V = I2 * R2
V = 3 * 40
V = 120v

(a)
Current in the first resistor -
V = I1 *R1
I1 = V/R1
I1 = 120/38
I1 = 3.16 A
Current in the first resistor I1 = 3.16 A

(b)
Power dissipated = V*I = 120 * (3 +3.16)
Total power supplied to the two resistors P = 739.2 Watt